OpenStudy (anonymous):

Hay got a word problem and need help here it is There are two baseball players together they had 374 hits Johnson had 10 more than Foree. How many hit did each player have? I do not just want the answer I want to know how to set it up so I can work it. please help me I have had this problem come up 3xs and I always get it wrong.

OpenStudy (tiffanymak1996):

let x be johnson and y be foree

OpenStudy (tiffanymak1996):

ok?

OpenStudy (tiffanymak1996):

are you here?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

y-x-10=374 is this the problem

OpenStudy (anonymous):

I do not know how to set up the problem to work it

OpenStudy (tiffanymak1996):

now, do you know the relationship between x and y?

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

Lost

OpenStudy (tiffanymak1996):

there are two equations in here: one is x+y=374

OpenStudy (tiffanymak1996):

the other one is: y=x-10

OpenStudy (anonymous):

Johnson is 10 +y=374

OpenStudy (tiffanymak1996):

johnson(x) plus foree(y) has 374 hits together, thus x+y=374 is this ok?

OpenStudy (anonymous):

okayyyyyyy

OpenStudy (anonymous):

could you set it up for me

OpenStudy (tiffanymak1996):

what you do is that you substitute y=x-10 in x+y=374

OpenStudy (tiffanymak1996):

thus: x+(x-10)=374

OpenStudy (tiffanymak1996):

do you get it?

OpenStudy (anonymous):

2x-10x=374 46.75

OpenStudy (tiffanymak1996):

be careful! x+(x-10) is 2x-10

OpenStudy (anonymous):

oh my bad I multiplied the 10 by x

OpenStudy (tiffanymak1996):

it's okay :)

OpenStudy (anonymous):

192

OpenStudy (anonymous):

that is what johnson made right?

OpenStudy (tiffanymak1996):

yes

OpenStudy (anonymous):

so that means that foree made 182

OpenStudy (tiffanymak1996):

yes

OpenStudy (anonymous):

ok got it thanks will write this formula down for future recall

OpenStudy (anonymous):

thanks so much

OpenStudy (tiffanymak1996):

just separate relationships between x and y, so that the picture would be clearer, and be careful when substituting.

OpenStudy (anonymous):

yes, you explained it so well now i got it

OpenStudy (tiffanymak1996):

gr8, and congrats!